OpenStudy (anonymous):

Determine whether the function is concave up and where it is concave down. Also find all inflection points. T(t) = 2t - t^3

OpenStudy (solomonzelman):

can you find the first derivative (for the first step) ?

OpenStudy (solomonzelman):

(just apply the power rule to each term, and you can use T'(t) as your notation for the left side)

OpenStudy (anonymous):

T'(t) = 2 - 3t^2

OpenStudy (anonymous):

then i have to set the function equals to zero?

OpenStudy (solomonzelman):

no, you have to find the second derivative, and only then set it to equal zero.

OpenStudy (anonymous):

T'(t) = -6t

OpenStudy (solomonzelman):

yes, but you meant, T''(t)=-6t

OpenStudy (anonymous):

-6t = 0 t = 0?

OpenStudy (anonymous):

since 0/-6 = 0

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ T\prime \prime(t)=-6t }\)

OpenStudy (solomonzelman):

yes, t=0 is (the only) inflection point.

OpenStudy (solomonzelman):

when the \(\normalsize\color{blue}{ T\prime \prime(t) }\) is greater than zero, for what values of \(\normalsize\color{blue}{ t }\)?

OpenStudy (anonymous):

idk...

OpenStudy (solomonzelman):

you have \(\normalsize\color{blue}{ T\prime \prime(t)=-6t }\) what happens when t is negative, what happens when t is positive?

OpenStudy (solomonzelman):

if \(\normalsize\color{blue}{ t }\) is a positive number, (starting from any value that is greater than zero), then the \(\normalsize\color{blue}{ T\prime \prime(t) }\) is less than zero. Right?

OpenStudy (anonymous):

ye

OpenStudy (solomonzelman):

What can you say about \(\normalsize\color{blue}{ T\prime \prime(t) }\) when \(\normalsize\color{blue}{ t }\) is a negative number?

OpenStudy (anonymous):

T''(t) is greater than zero

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ T\prime \prime(t)=-6t }\) So: If, \(\normalsize\color{blue}{ t>0 }\) then, \(\normalsize\color{blue}{ T\prime \prime(t)<0 }\) AND If, \(\normalsize\color{blue}{ t<0 }\) then, \(\normalsize\color{blue}{ T\prime \prime(t)>0 }\)

OpenStudy (solomonzelman):

can you give me the concavity in interval notation?

OpenStudy (anonymous):

concave down and concave up?

OpenStudy (solomonzelman):

yes, one interval for concave up, and the other is concave down. (I have even asked the question in the needed for you order :) )

OpenStudy (solomonzelman):

I'll give you two intervals, and you will tell me which is concave up, and which is concave down, okay?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

1) \(\normalsize\color{blue}{ (-\infty,0) }\) 2) \(\normalsize\color{blue}{ (0,+\infty) }\)

OpenStudy (anonymous):

1. concave up 2. concave up?

OpenStudy (solomonzelman):

can't be that both are concave up.

OpenStudy (anonymous):

T''(t) > 0 = concave up T''(t) < 0 = concave down

OpenStudy (solomonzelman):

yes, but on what interval is T''(t)>0 and on what interval is T''(t)<0 ? (reminding that T''(t)=-6t in our case)

OpenStudy (anonymous):

\[(-\infty, 0) concave down ||(0,+\infty) concave up\]

OpenStudy (solomonzelman):

it is the other way, and also you can use ~ for a space.

OpenStudy (anonymous):

how would i sketch the graph for this?

OpenStudy (solomonzelman):

I wouldn't just sketch it based on the concavity. although you can make a sketch something like